A compound contains 46.7% N and 53.3% O. If the molecular mass of the compound is 60.0g/mol. what is the empirical formula?
1
Expert's answer
2021-10-24T01:28:11-0400

46.7% = Nitrogen

53.3% = Oxygen


Step1: Divide by their molar mass es

46.7/14 = 3.34 ​Nitrogen

53.3/16 = 3.33 ​Oxygen


Step2: Divide by the lowest result

3.34/3.33 ≈1 ​Nitrogen

3.33/3.33 ≈ 1 ​Oxygen


The empirical formula is therefore NO


The molecular mass is 60.0 g/mol

∴(NO)n=60

(14+16)n=60

(30)n=60

n=2


∴ ​The molecular formula is N2O2"."